OpenStudy (anonymous):

Write the expression as a single logarithm. \[5\log_{b}y + 4\log_{b}x\]

OpenStudy (anonymous):

\[\log_{b}y^5 + \log_{b}x^4\] Property: \[ \log A +\log B = \log AB\] \[\huge\log_{b}y^5x^4\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

@Luigi0210 sir can you check it I did this correctly?

OpenStudy (luigi0210):

That is correct :)

OpenStudy (anonymous):

Thanks :)